OpenStudy (anonymous):

What is the sum of the roots of the polynomial shown below? F(x)=3x^3+12x^2+3x-18

OpenStudy (anonymous):

F(x)=(x-1)(3x^2+15x+18)=3(x-1)(x+2)(x+3) roots: x=1,-2,-3 sum of these is -4 (why do you want the sum?)

OpenStudy (anonymous):

Thats what the question says... (the sum)

OpenStudy (anonymous):

ax^3+bx^2+cx+d=0 roots x1, x2, x3 then x1+x2+x3=-b/a so x1+x2+x3=-12/3=-4

OpenStudy (anonymous):

sum of roots=-(12/3)=-4